# Database Views as Explanations for Relational Deep Learning

### v1

@@ 3. RDL Explanations via Soft Determinacy
In this section, we introduce our framework, where we leverage the concept of view determinacy to produce explanations.
@@ 3.1. Explanation Views
create view $V\mathord{\langle s\rangle}$ as
select * from $T$
where $T.A=s.A$
Note that the parameterized view $V$ does not necessarily have to refer to its parameter in its definition. If $V$ ignores its parameter $s$, we say that it is instance-agnostic, otherwise it is instance-specific. Formally, $V$ is instance-agnostic if, for each database $D$ and instances $s$ and $s^{\prime}$ it holds that $V\mathord{\langle s\rangle}(D)=V\mathord{\langle s^{\prime}\rangle}(D)$. As an example, the following explanation view is instance-agnostic, as it refers to all tuples of $T$, regardless of their relationship to $s$:
select * from $T$
Finally, by an explanation $E$ we mean a finite subset of concrete queries from $\mathbf{EL}$. The quality of an explanation is measured using two criteria: determinacy (i.e., whether $E$ is enough to determine the behavior of $M$) and conciseness (i.e., how easy it is to grasp $E$).
@@ 3.2. Soft Determinacy
Consider an explanation $E$ of an $R$-model $M$ on a database $D$. Intuitively, $E$ “sufficiently determines” $M$ if $M$ behaves similarly even if $D$ changes, as long as the result of $E$ is not affected by this change. Ideally, for any two databases that agree on the explanation views in $E$, the output of model $M$ should be identical. Viewing $M$ as a query, this strict sufficiency notion is precisely determinacy, as described in Section 2.1. To capture the imperfect nature of explanations more effectively, we relax the notion of determinacy in two dimensions. First, we do not require exact determinacy; instead, we account for approximate agreement by quantifying the distance between model predictions before and after the change (see Figure 1). Second, we move away from deterministic determinacy by providing a probabilistic adaptation. For that, we require determinacy in expectation, using our original database $D$ as reference and assuming a probability distribution of contingency databases. Applying the two relaxations leads to an approximate probabilistic notion of determinacy, namely soft determinacy.
Formally, we assume a distribution $\Delta$ over contingency databases that agree on the schema $\mathcal{S}$ of $D$. To be comparable to $M(D)$, we require each database $D^{\prime}$ in this distribution to agree with $D$ on the keys of $R$, that is:
$D^{\prime}(R)[\mathsf{key}(R)]=D(R)[\mathsf{key}(R)]\,.$
In order to define approximate agreement, suppose that $M$ maps an instance $s$ to $a$, and an instance $s^{\prime}$ ito $a^{\prime}$. We use a distance metric dist that gives a quantity $\textrm{dist}(a,a^{\prime})$ of how far $a$ is from $a^{\prime}$. For example, consider the special case where $M$ is a binary classifier with predictions $\in\{0,1\}$; then, it makes sense to use $d$ that checks for identity: $\textrm{dist}(a,a^{\prime})$ is 0 if $a=a^{\prime}$, and 1 otherwise.
Using the above, we arrive at the property of soft determinacy for an explanation $E$, quantified by deviation from determinacy, so that lower deviation implies a better explanation.
@@
Deviation from determinacy of $E$ is then the average over all instances:
$\mathsf{dev}_{\Delta}(E)\vcentcolon=\frac{1}{|D(R)|}\sum_{s\in D(R)[\mathsf{key}(R)]}\hskip-11.38109pt\mathsf{dev}_{\Delta}(E,s).$
@@ 3.3. Explanation Conciseness
We associate with each $V\in\mathbf{EL}$ a cost related to its complexity. This cost can be syntactic, e.g., the number of symbols involved in the SQL representation. Alternatively, it can be semantic, e.g., the number of tuples in the view $|V(D)|$. We denote this cost by $\mathsf{cost}(V)$. We then define the cost of an explanation $E$ as $\mathsf{cost}(E)=\sum_{V\in E}\mathsf{cost}(V)$. Note that the syntactic cost does not depend on the database $D$, as the SQL representation refers only to the schema $\mathcal{S}$.
@@
for some trade-off captured by $\lambda\in\mathbb{R},\lambda\geq 0$.
Clearly, there is a natural trade-off between achieving a high level of determinacy (i.e., the minimization of $\mathsf{dev}_{\Delta}(E)$) and low explanation complexity (the minimization of $\mathsf{cost}(E)$). Consider the extreme case of $E$ containing the entire schema $\mathcal{S}$. In this case, $E$ fully determines the model, but also has the complexity of the model that we wish to explain to begin with. On the other hand, a completely uninformative explanation $E=\emptyset$ has minimal complexity, but is likely to hide what the model requires for its predictions.
The computational objective involves the estimation of a conditional probability, a generally intractable task, in order to compute $\mathsf{dev}_{\Delta}(E)$. In Section 4 we discuss several choices for realizing this objective into a tractable optimization problem. Additionally, in Section 5 we propose a learning-based solution for efficient explanation view discovery in the case of GNNs.
@@ 4.1. Explanation Languages
These views retain all tuples produced by joining $R$ and $S$ using foreign-key references. Note that more complex join patterns can be captured by multiple views in an explanation view set. Because FKJoin focuses on foreign-key attributes, it always preserves all key attributes as well as the data attributes of relations that appear in at least one explanation view.
@@
where $\phi$ is a selection predicate that identifies an informative subset of tuples. We assume that Selection does not affect the graph topology, hence we implicitly assume that an explanation in Selection includes all views of the form “select $\mathsf{key}(R)$, $\mathord{\mathbf{F}}$ from RR” for every $R\in\mathsf{rel}(\mathcal{S})$, where $\mathsf{key}(R)$ contains all key attributes and $\mathord{\mathbf{F}}$ contains all foreign-key attributes of $\mathcal{S}(R)$.
For example, consider a database with a relation Product $P$. The explanation view
select * from $P$ where $P.price<50$ and $P.cat=$ “electronics”
indicates that low-priced electronics constitute an informative region of the data for the model.
Cost model
Explanation conciseness requires a precise definition of the cost model so that every explanation view $V\in E$ is associated with a value $\mathsf{cost}(V)$. Here we propose a simple cost model that focuses on minimizing the description cost of explanation views, making them easier for a user to understand. For Projection, $\mathsf{cost}(V)$ is the number of highlighted data attributes, i.e., $\mathsf{cost}(V)=k$. For FKJoin, each binary join incurs a cost of $1$, i.e., $\mathsf{cost}(V)=1$. For Selection, $\mathsf{cost}(V)$ depends on the complexity of the selection predicate $\phi$. Later, we will consider disjunctions of the form $\phi_{1}$ or $\dots$ or $\phi_{l}$, where each $\phi_{i}$ is an atomic predicate. In this case we simply assign $\mathsf{cost}(V)=l$, i.e., the number of atomic predicates involved.
Language composition
We consider combinations of the three languages by constructing composite explanation views in the relational algebra sense. Importantly, these combination languages produce composite views rather than unions of separate explanation views belonging to one of the three defined languages. For composite views, the cost model is naturally extended: we sum the number of data attributes in projections, the number of joins, and the number of selection predicates involved.
@@ 4.2. Database Perturbations
Let us consider two extreme scenarios to illustrate the impact of the choice of $\Delta$. First, consider a $\Delta$ over databases $D^{\prime}$ that differ from $D$ in just one value of a single tuple. A well-behaved model would likely give a similar output on $D$ and $D^{\prime}$, irrespective of the chosen explanation $E$. Therefore, such $\Delta$ would not yield a discriminative criterion between candidate explanations. On the other hand, suppose $\Delta$ is the uniform distribution over all possible $D^{\prime}$, no matter how different they are from $D$. In this case, it is not clear how one could obtain samples from $\Delta$ and, more importantly, the sampled $D^{\prime}$ could be unrepresentative for the application domain of the original database $D$. Ideally, we would like to limit the explanatory power of $E$ to the useful regime, i.e., restricting $\Delta$ to a local neighborhood of $D$ that captures realistic variations while preserving relevance to the application domain.
For each explanation language $\mathbf{EL}$, we give a number of strategies for realizing $\Delta$ by perturbing the original database $D$. The resulting databases $D^{\prime}$ are substantially different from $D$, yet maintain certain properties of $D$, achieving a balance in the aforementioned tradeoff.
(1) Projection
The core idea is to permute the values of columns that are not used in any projection explanation, inspired by a well-known model-agnostic feature selection technique, namely permutation feature importance (PFI) (Breiman, 2001; Fisher et al., 2019; Molnar, 2025). Such column-wise permutations respect the active domain of each attribute and preserve aggregate statistics within a column, while breaking correlations between a permuted column and the model output.
@@
Formally, let $\mathsf{Attr}(E)$ be the set of attributes used in views of projection explanations $E$. Recall that $\mathsf{Attr}(E)$ includes all key attributes. Because explanation views must be respected by $\Delta$, we can only permute attributes not present in projection explanations. For each relation name $R\in\mathsf{rel}(\mathcal{S})$, let $\mathsf{attPerm}(R)$ be the set of attributes in $\mathsf{att}(R)$ not in $\mathsf{Attr}(E)$, which are allowed to be permuted.
Let $r=D(R)$. A relation $r^{\prime}$ is a column-wise permutation of $r$ under Projection explanation $E$, written $r^{\prime}\in\mathsf{perm}_{E}(r)$, if $r^{\prime}$ has the same schema and cardinality as $r$, and for each tuple $t^{\prime}\in r^{\prime}$ there exists a unique tuple $t\in r$ with $t^{\prime}[\mathsf{key}(R)]=t[\mathsf{key}(R)]$ such that for all $A\in\mathsf{attPerm}(R)$, $t^{\prime}[A]$ is a permutation of the multiset of values of the attribute $A$ in $r$, while for $A\notin\mathsf{attPerm}(R)$, $t^{\prime}[A]=t[A]$. The joint column-wise permutation $r^{\prime}$ of $r$, written $r^{\prime}\in\mathsf{jointperm}_{E}(r)$, is defined similarly, but all attributes in $\mathsf{attPerm}(R)$ are permuted by the same permutation, preserving $t^{\prime}[\mathsf{key}(R)]=t[\mathsf{key}(R)]$ for the unique corresponding $t\in r$.
The column-wise permutation distribution $\Delta^{\mathsf{ind}}_{E}$ is the uniform distribution with support: $\{D^{\prime}\mid\forall R\in\mathsf{rel}(\mathcal{S}):D^{\prime}(R)\in\mathsf{perm}_{E}(D(R))\}$. The joint column-wise permutation distribution $\Delta^{\mathsf{joint}}_{E}$ is the uniform distribution with support: $\{D^{\prime}\mid\forall R\in\mathsf{rel}(\mathcal{S}):D^{\prime}(R)\in\mathsf{jointperm}_{E}(D(R))\}$. In Section 6 we conduct experiments with both $\Delta^{\mathsf{ind}}_{E}$ and $\Delta^{\mathsf{joint}}_{E}$ and show that the deviation from determinacy is robust to this choice.
@@
(2) FKJoin
In the case of FKJoin, the goal of a perturbation is to make joins uninformative unless they appear in $E$. Since the joins we consider always correspond foreign-key constraints, we apply a perturbation on the foreign key side. This perturbs the target join without affecting other joins on the side of the primary key.
@@
The example in Figure 3 (left) shows perturbations applied to the primary-foreign pair $T_{2}.B$ - $T_{1}.B$. One could consider permuting the foreign key $T_{1}.B$ as in a projection explanation. But this strategy would not be effective in breaking the predictive signal encoded in the actual join condition as it would retain the number of primary-foreign key pairs per unique value of $T_{1}.B$, i.e., $1$ would appear 4 times while $2$ and $3$ would each appear once. In order to break the join information, we randomly replace each value of $T_{1}.B$ with a value in the domain $\{1,2,3\}$, either uniformly at random (right) or mimicking the underlying frequency distribution of foreign keys (middle). In the former each value in $\{1,2,3\}$ appears $2$ times while in the latter $2$ appears 4 times while $1$ and $3$ each appear once, preserving the frequencies $4,1,1$.
Formally, let $\mathsf{FK}(E)$ denote the set of foreign keys that appear in the explanation $E$. For each $R\in\mathsf{rel}(\mathcal{S})$ and $r=D(R)$, we define a perturbed relation $r^{\prime}$ by independently replacing the values of foreign-key attributes outside of $\mathsf{FK}(E)$ with other values from their respective domains. We consider two alternative strategies for generating $r^{\prime}$. In the uniform replacement $\Delta^{\mathsf{uniform}}_{E}$, each foreign key value is replaced independently and uniformly at random from its domain. In the frequency-preserving replacement $\Delta^{\mathsf{freq}}_{E}$, the multiset of foreign key values in the perturbed relation matches the multiset of frequencies in the original relation: if a foreign key value occurs $i$ times in $r$, then there exists some foreign key value in $r^{\prime}$ that occurs $i$ times, although not necessarily in the same tuples.
Figure 3. Example 2: foreign key perturbations for FKJoin. Here, the foreign key $T_{1}.B$ is perturbed as it is is not in $\mathsf{FK}(E)$.
(3) Selection
For Selection we adapt the column-wise permutation strategies used in the case of Projection. Here, we permute all data attributes only for the tuples that do not satisfy the selection condition of the explanation view. We provide an example in Figure 2 (lower). Specifically, let $\mathsf{Tups}(E)\subseteq\bigcup\{D(R)|R\in\mathsf{rel}(\mathcal{S})\}$ be the set of tuples that satisfy the selection condition. Each relation $r=D(R)$ is separated into $r_{E}$ and $r_{E}^{\prime}$ such that $r_{E}\subseteq\mathsf{Tups}(E)$ and $r_{E}^{\prime}=r\setminus r_{E}$. We then apply column-wise permutations, either independently or jointly, only in $r_{E}^{\prime}$ with $\mathsf{Attr}(E)$ including all data attributes, i.e., only preserving key and foreign key attributes.

### v2

@@ 3. RDL Explanations via Soft Determinacy
In this section, we introduce our framework, where we leverage the concept of determinacy to produce explanations.
@@ 3.1. Explanation Views
create view $V\mathord{\langle s\rangle}$ as select * from $T$ where $T.A=s.A$
Note that the parameterized view $V$ does not necessarily have to refer to its parameter in its definition. If $V$ ignores its parameter $s$, we say that it is instance-agnostic, otherwise it is instance-specific. Formally, $V$ is instance-agnostic if, for each database $D$ and instances $s$ and $s^{\prime}$ it holds that $V\mathord{\langle s\rangle}(D)=V\mathord{\langle s^{\prime}\rangle}(D)$. As an example, the following explanation view is instance-agnostic, as it refers to all tuples of $T$, regardless of their relationship to $s$: select * from $T$.Finally, by an explanation $E$ we mean a finite subset of concrete queries from $\mathbf{EL}$. The quality of an explanation is measured using two criteria: determinacy (i.e., whether $E$ is enough to determine the behavior of $M$) and conciseness (i.e., how easy it is to grasp $E$).
@@ 3.2. Soft Determinacy
Consider an explanation $E$ of an $R$-model $M$ on a database $D$. Intuitively, $E$ “sufficiently determines” $M$ if $M$ behaves similarly when $D$ changes, as long as the result of $E$ is not affected by this change. Ideally, for any two databases that agree on the explanation views in $E$, the output of $M$ should be identical. Viewing $M$ as a query, this strict sufficiency notion is precisely determinacy, as described in Section 2.1. To capture the imperfect nature of explanations , we relax the notion of determinacy in two dimensions. First, we do not require exact determinacy; instead, we account for approximate agreement by quantifying the distance between model predictions before and after the change (see Figure 1). Second, we move away from deterministic determinacy via a probabilistic adaptation. For that, we require determinacy in expectation, using our original database $D$ as reference and assuming a probability distribution of contingency databases. The two relaxations lead to an approximate probabilistic notion of determinacy, namely soft determinacy.
Formally, we assume a distribution $\Delta$ over contingency databases that agree on the schema $\mathcal{S}$ of $D$. To be comparable to $M(D)$, we require each database $D^{\prime}$ in this distribution to agree with $D$ on the keys of $R$, that is: $D^{\prime}(R)[\mathsf{key}(R)]=D(R)[\mathsf{key}(R)]\,$.
To define approximate agreement, suppose that $M$ maps an instance $s$ to $a$, and an instance $s^{\prime}$ to $a^{\prime}$. We use a distance metric dist that quantifies how far $a$ is from $a^{\prime}$ as $\textrm{dist}(a,a^{\prime})$. For example, consider the special case where $M$ is a binary classifier with predictions $\in\{0,1\}$; then, it makes sense to use $d$ that checks for identity: $\textrm{dist}(a,a^{\prime})$ is 0 if $a=a^{\prime}$, and 1 otherwise.
Using the above, we arrive at the soft determinacy property for an explanation $E$, quantified by deviation from determinacy, so that lower deviation implies a better explanation.
@@
Deviation from determinacy of $E$ is the empirical mean over instances: $\mathsf{dev}_{\Delta}(E)\vcentcolon=\frac{1}{|D(R)|}\sum_{s\in D(R)[\mathsf{key}(R)]}\mathsf{dev}_{\Delta}(E,s)$.
@@ 3.3. Explanation Conciseness
We associate with each $V\in\mathbf{EL}$ a cost related to its complexity. This cost can be syntactic, e.g., the number of symbols involved in the SQL representation. It can also be semantic, e.g., the number of tuples in the view $|V(D)|$. We denote this cost by $\mathsf{cost}(V)$. We then define the cost of an explanation $E$ as $\mathsf{cost}(E)=\sum_{V\in E}\mathsf{cost}(V)$.
@@
for a trade-off captured by $\lambda\in\mathbb{R},\lambda\geq 0$.
Clearly, there is a natural trade-off between achieving high soft determinacy (minimization of $\mathsf{dev}_{\Delta}(E)$) and low explanation complexity (minimization of $\mathsf{cost}(E)$). Consider the extreme case of $E$ containing the entire schema $\mathcal{S}$. In this case, $E$ fully determines the model $M$, but also does not reduce the complexity of $M$ or $\mathcal{S}$. On the other hand, a void explanation $E=\emptyset$ has minimal complexity, but likely does not capture any insight on the model’s behavior.
The optimization problem in Definition 3 displays the challenges commonly associated with bilevel optimization (Dempe, 2002; Colson et al., 2007). It involves two levels: the outer selects explanation views to minimize cost, while the inner estimates $\mathsf{dev}_{\Delta}(E)$ by sampling and evaluating the model on multiple perturbed databases that respect the constraints imposed by the outer explanation selection. The outer level, i.e., searching among exponentially-many subsets of database components to find those that best preserve the model behavior, requires exploring a vast combinatorial space (see Example 4).
Example 4.
Consider the case where we seek explanation views that only involve projections over a single table with binary attributes, and we require exact determinacy ($\mathsf{dev}_{\Delta}(E)=0$). Suppose our model $M$ is a disjunction (OR) over all attributes, i.e., $M$ predicts $1$ for all tuples except the all-zeros tuple $(0,\ldots,0)$. A minimal explanation corresponds to a minimal set of columns such that every tuple has at least one $1$ in those columns. This is precisely the classical hitting set problem, which is NP-hard. This example demonstrates that even for the simplest explanation language and model structure, finding optimal explanations requires considering exponentially many column subsets. The problem becomes more complex with more expressive explanations not restricted to projections.
To address these challenges, in Section 4 we discuss several choices for realizing this objective into a more practical optimization problem. Additionally, in Section 5 we propose a learning-based solution for efficient explanation view discovery for hetero-GNNs.
@@ 4.1. Explanation Languages ($\mathbf{EL}$’s)
These views retain all tuples produced by joining $R$ and $S$ using foreign-key references. Note that more complex join patterns can be captured by a set of multiple views. Because FKJoin focuses on foreign keys, it always preserves all key attributes and data attributes of relations that appear in at least one explanation view.
@@
where $\phi$ is a selection predicate that identifies an informative subset of tuples. For example, consider a database with a Product $P$ relation. The explanation view: “select * from $P$ where $P.price<50$ and $P.cat=$ Electronics” highlights the group of affordable electronics.
Cost model. Explanation conciseness requires a precise definition of the cost model so that every explanation view $V\in E$ is associated with a value $\mathsf{cost}(V)$. Here we propose a simple cost model that focuses on minimizing the description cost of explanation views, making them easier for a user to understand. For Projection, $\mathsf{cost}(V)$ is the number of highlighted data attributes, i.e., $\mathsf{cost}(V)=k$. For FKJoin, each binary join incurs a cost of $1$, i.e., $\mathsf{cost}(V)=1$. For Selection, $\mathsf{cost}(V)$ depends on the complexity of the selection predicate $\phi$. Later, we will consider disjunctions of the form $\phi_{1}$ or $\dots$ or $\phi_{l}$, where each $\phi_{i}$ is an atomic predicate. In this case we simply assign $\mathsf{cost}(V)=l$, i.e., the number of atomic predicates involved.
Language composition. We consider combinations of the three languages by constructing composite explanation views in the relational algebra sense. Importantly, these combination languages produce composite views rather than unions of separate explanation views belonging to one of the three defined languages. For composite views, the cost model is naturally extended: we sum the number of data attributes in projections, the number of joins, and the number of selection predicates involved.
@@ 4.2. Database Perturbations
Let us consider two extreme scenarios to illustrate the impact of $\Delta$. First, consider a $\Delta$ over databases $D^{\prime}$ that differ from $D$ in just one value of a single tuple. A well-behaved model would likely give a similar output on $D$ and $D^{\prime}$, irrespective of the chosen explanation $E$. Therefore, such $\Delta$ would not yield a discriminative criterion between candidate explanations. On the other hand, if $\Delta$ is the uniform distribution over all possible $D^{\prime}$, no matter how different they are from $D$, it is not clear how one could obtain samples from $\Delta$. More importantly, the sampled $D^{\prime}$ could be unrepresentative for the application domain of the original database $D$. Ideally, we would like to limit the explanatory power of $E$ to the useful regime, i.e., restricting $\Delta$ to a local neighborhood of $D$ that captures realistic variations while preserving relevance to the application domain.
Defining input perturbations is a fundamental challenge across ML explanation approaches that replace insignificant features with uninformative values. Permutation Feature Importance (PFI), a well-known model-agnostic feature selection technique (Breiman, 2001; Molnar, 2025), replaces feature values by permuting them across instances, preserving marginal statistics while breaking correlations with the target. However, this approach may generate unrealistic data variations that violate feature correlations, thus forcing models to extrapolate (Fisher et al., 2019; Hooker et al., 2021). PFI variants address this by sampling from the conditional distribution $P(x_{j}|x_{-j})$ of the perturbed feature $x_{j}$ (Zintgraf et al., 2017; Fisher et al., 2019; Debeer and Strobl, 2020), but become computationally inefficient without strong assumptions like feature locality (Zintgraf et al., 2017), assumptions more natural for images than tabular data. Subgroup-based methods partition instances into groups in order to reduce dependence before permuting (Molnar et al., 2023), however they introduce additional complexity by training auxiliary models (e.g., decision trees) to identify appropriate subgroups for each feature.
We adapt these insights to relational databases by designing language-specific perturbations. For each explanation language $\mathbf{EL}$, we provide strategies for realizing $\Delta$ by perturbing the original database $D$. Each strategy neutralizes information outside the explanation while respecting the explanation views. Following the PFI literature’s concern with correlation preservation (Fisher et al., 2019; Hooker et al., 2021; Molnar et al., 2023), we test both correlation-preserving and correlation-breaking variants in Section 6 and find results are mostly robust to this choice.
Projection. We permute columns not appearing in the explanation (independently or jointly), inspired by PFI.
@@
Formally, let $\mathsf{Attr}(E)$ be the set of attributes used in a projection explanation $E$. Recall that $\mathsf{Attr}(E)$ includes all key attributes. Because explanation views must be respected by $\Delta$, we only permute attributes not in $\mathsf{Attr}(E)$. We define two permutation distributions that differ in how they handle correlations among non important attributes. The joint column-wise permutation distribution $\Delta^{\mathsf{joint}}_{E}$ is the uniform distribution over all databases $D^{\prime}$ where, for each relation in $D$, attributes $\notin\mathsf{Attr}(E)$ are permuted together by the same permutation. This preserves correlations among unimportant attributes while destroying correlations between important and unimportant attributes, and between the target and unimportant attributes. The independent column-wise permutation distribution $\Delta^{\mathsf{ind}}_{E}$ is the uniform distribution over all databases $D^{\prime}$ where each attribute $\notin\mathsf{Attr}(E)$ is permuted independently. This additionally breaks correlations among unimportant attributes.
@@
FKJoin. The goal of a perturbation is to make joins uninformative unless they appear in $E$. Since the joins we consider always correspond foreign-key constraints, we apply a perturbation on the foreign key side. This perturbs the target join without affecting other joins on the side of the primary key.
@@
Figure 3 (left) shows perturbations applied to the primary-foreign pair $T_{2}.B$ - $T_{1}.B$. One could consider permuting the foreign key $T_{1}.B$. But this strategy would not be effective in breaking the predictive signal encoded in the actual join condition as it would retain the number of primary-foreign key pairs per unique value of $T_{1}.B$, i.e., $1$ would appear 4 times while $2$ and $3$ would each appear once. In order to break the join information, we randomly replace each value of $T_{1}.B$ with a value in the domain $\{1,2,3\}$, either uniformly at random (right) or mimicking the underlying frequency distribution of foreign keys (middle). In the former each value in $\{1,2,3\}$ appears $2$ times while in the latter $2$ appears 4 times while $1$ and $3$ each appear once, preserving the frequencies $4,1,1$.
We consider two alternative strategies: In uniform replacement $\Delta^{\mathsf{uniform}}_{E}$, each foreign key value is replaced independently and uniformly at random from its domain. In the frequency-preserving replacement $\Delta^{\mathsf{freq}}_{E}$, the frequences of foreign key values in the original relation $r$ and the perturbed one $r^{\prime}$ match: if a value occurs $i$ times in $r$, then there exists some value in $r^{\prime}$ that occurs $i$ times, although not necessarily in the same tuples. Figure 3 shows an example, with $\mathsf{FK}(E)$ denoting the foreign keys in explanation $E$.
Figure 3. Example 2: foreign key perturbations for FKJoin. Here, the foreign key $T_{1}.B$ is perturbed as it is not in $\mathsf{FK}(E)$.
Selection. We adapt the column-wise permutation strategies used in the case of Projection for all data attributes, but applied only to the tuples that do not satisfy the selection condition of the explanation view. Figure 2 (lower) shows an example, with $\mathsf{Tups}(E)$ denoting the set of tuples that satisfy the selection condition.